# Scalable MatMul-free Language Modeling

### v5

@@ Abstract
Matrix multiplication (MatMul) typically dominates the overall computational cost of large language models (LLMs). This cost only grows as LLMs scale to larger embedding dimensions and context lengths. In this work, we show that MatMul operations can be completely eliminated from LLMs while maintaining strong performance at billion-parameter scales. Our experiments show that our proposed MatMul-free models achieve performance on-par with state-of-the-art Transformers that require far more memory during inference at a scale up to at least 2.7B parameters. We investigate the scaling laws and find that the performance gap between our MatMul-free models and full precision Transformers narrows as the model size increases. We also provide a GPU-efficient implementation of this model which reduces memory usage by up to 61% over an unoptimized baseline during training. By utilizing an optimized kernel during inference, our model’s memory consumption can be reduced by more than 10$\times$ compared to unoptimized models. To properly quantify the efficiency of our architecture, we build a custom hardware solution on an FPGA which exploits lightweight operations beyond what GPUs are capable of. We processed billion-parameter scale models at 13W beyond human readable throughput, moving LLMs closer to brain-like efficiency. This work not only shows how far LLMs can be stripped back while still performing effectively, but also points at the types of operations future accelerators should be optimized for in processing the next generation of lightweight LLMs. Our code implementation is available at https://github.com/ridgerchu/matmulfreellm.
1 Introduction
@@
Matrix Multiplication (MatMul) is the dominant operation in most neural networks, where dense layers involve vector-matrix multiplication (VMM), convolutions can be implemented as block-sparse VMMs with shared weights, and self-attention relies on matrix-matrix multiplication (MMM). The prevalence of MatMul is primarily due to Graphics Processing Units (GPUs) being optimized for MatMul operations. By leveraging Compute Unified Device Architecture (CUDA) and highly optimized linear algebra libraries such as cuBLAS, the MatMul operation can be efficiently parallelized and accelerated. This optimization was a key factor in the victory of AlexNet in the ILSVRC2012 competition and a historic marker for the rise of deep learning [1]. AlexNet notably utilized GPUs to boost training speed beyond CPU capabilities, and as such, deep learning won the ‘hardware lottery’ [2]. It also helped that both training and inference rely on MatMul.
@@
Despite its prevalence in deep learning, MatMul operations account for the dominant portion of computational expense, often consuming the majority of the execution time and memory access during both training and inference phases. Several works have replaced MatMul with simpler operations through two main strategies. The first strategy involves substituting MatMul with elementary operations, e.g., AdderNet replaces multiplication with signed addition in convolutional neural networks (CNNs) [3]. Given the focus on convolutions, AdderNet is intended for use in computer vision over language modeling.
@@
The second approach employs binary or ternary quantization, simplifying MatMul to operations where values are either flipped or zeroed out before accumulation. Quantization can be applied to either activations or weights: spiking neural networks (SNNs) use binarized activations [4, 5, 6], while binary and ternary neural networks (BNNs and TNNs) use quantized weights [7]. Both methods can also be combined [8, 9].
@@
Recent advances in language modeling, like BitNet [10, 11], demonstrate quantization’s scalability, replacing all dense layer weights with binary/ternary values to support up to 3 billion parameters. Despite replacing VMMs with accumulations in all dense layers, BitNet retains the self-attention mechanism which relies on an expensive MMM. Dynamically computed matrices $Q$ (query) and $K$ (key) are multiplied to form the attention map. Since both $Q$ and $K$ matrices are dynamically computed from pre-activation values, achieving optimal hardware efficiency on GPUs requires custom optimizations, such as specialized kernels and advanced memory access patterns. Despite these efforts, such MatMul operations remain resource-intensive on GPUs, as they involve extensive data movement and synchronization which can significantly hinder computational throughput and efficiency [12]. In our experiments, ternary quantization of the attention matrices in BitNet causes a significant drop in performance and failure to reach model convergence (see Fig. 1). This raises the question: is it possible to completely eliminate MatMul from LLMs?
@@
In this work, we develop the first scalable MatMul-free language model (Matmul-free LM) by using additive operations in dense layers and element-wise Hadamard products for self-attention-like functions. Specifically, ternary weights eliminate MatMul in dense layers, similar to BNNs. To remove MatMul from self-attention, we optimize the Gated Recurrent Unit (GRU) [13] to rely solely on element-wise products and show that this model competes with state-of-the-art Transformers while eliminating all MatMul operations.
To quantify the hardware benefits of lightweight models, we provide an optimized GPU implementation in addition to a custom FPGA accelerator. By using fused kernels in the GPU implementation of the ternary dense layers, training is accelerated by 25.6% and memory consumption is reduced by up to 61.0% over an unoptimized baseline on GPU. Furthermore, by employing lower-bit optimized CUDA kernels, inference speed is increased by 4.57 times, and memory usage is reduced by a factor of 10 when the model is scaled up to 13B parameters. This work goes beyond software-only implementations of lightweight models and shows how scalable, yet lightweight, language models can both reduce computational demands and energy use in the real-world.
2 Related Works
Binary, Ternary, and Low-Precision Quantization for Language Models:
The effort to quantize language models began with reducing a ternary BERT into a binarized model [14], achieving 41% average accuracy on the GLUE benchmarks with subsequent fine-tuning. Ref. [15] distilled the intermediate outputs from a full precision BERT to a quantized version. Recently, Ref. [16] introduced an incremental quantization approach, progressively quantizing a model from 32-bit to 4-bit, 2-bit, and finally to binary model parameters. Following the quantization of BERT, low-precision language generation models have gained momentum. Ref. [17] used Quantization-Aware Training (QAT) to successfully train a model with 2-bit weights. BitNet pushed this to 3-billion-parameter binary and ternary models while maintaining competitive performance with Llama-like language models [10, 11].
MatMul-free Transformers:
The use of MatMul-free Transformers has been largely concentrated in the domain of SNNs. Spikformer led the first integration of the Transformer architecture with SNNs [18, 19], with later work developing alternative Spike-driven Transformers [20, 21]. These techniques demonstrated success in vision tasks. In the language understanding domain, SpikingBERT [22] and SpikeBERT [23] applied SNNs to BERT utilizing knowledge distillation techniques to perform sentiment analysis. In language generation, SpikeGPT trained a 216M-parameter generative model using a spiking RWKV architecture. However, these models remain constrained in size, with SpikeGPT being the largest, reflecting the challenges of scaling with binarized activations. In addition to SNNs, BNNs have also made significant progress in this area. BinaryViT [24] and BiViT [25] successfully applied Binary Vision Transformers to visual tasks. Beyond these approaches, Kosson et al. [26] achieve multiplication-free training by replacing multiplications, divisions, and non-linearities with piecewise affine approximations while maintaining performance.
3 Method
In this section, we break down the components of the proposed MatMul-free LM. We first describe the MatMul-free dense layers (BitLinear layers) that use ternary weights. By constraining the weights to the set $\{-1,0,+1\}$ and applying additional quantization techniques, MatMul operations are replaced with addition and negation operations. This reduces computational cost and memory utilization, while preserving the expressiveness of the network. We then provide further detail of our MatMul-free LM architecture, which includes a token mixer for capturing sequential dependencies and a channel mixer for integrating information across embedding dimensions.
The Method section is structured as follows. First, in Sec. 3.1, we provide a comprehensive description of the MatMul-free dense layers with ternary weights, which form the foundation of our approach. Next, Sec. 3.2 introduces our hardware-efficient fused BitLinear layer, designed to optimize the implementation of BitLinear layers. Building upon these components, Sec. 3.3 delves into the details of our MatMul-free LM architecture. We present the MatMul-free token mixer, where we propose the MatMul-free Linear Gated Recurrent Unit (MLGRU), and the MatMul-free channel mixer, which employs the Gated Linear Unit (GLU) with BitLinear layers. By combining the MLGRU token mixer and the GLU channel mixer with ternary weights, our proposed architecture relies solely on addition and element-wise products. Finally, Sec. 3.4 provides an overview of the training details used to optimize our model.
@@ 3.1 MatMul-free Dense Layers with Ternary Weights
In a standard dense layer, the MatMul between the input $x\in\mathbb{R}^{1\times d}$ and the weight matrix $W\in\mathbb{R}^{d\times m}$ can be expressed as:
@@
where $y\in\mathbb{R}^{1\times m}$ is the output. To avoid using standard MatMul-based dense layers, we adopt BitNet to replace dense layers containing MatMuls with BitLinear modules, which use ternary weights to transform MatMul operations into pure addition operation with accumulation, i.e., ternary accumulation. When using ternary weights, the elements from the weight matrix $W$ are constrained to values from the set $\{-1,0,+1\}$. Let $\widetilde{\mathbf{W}}$ denote the ternary weight matrix. The MatMul with ternary weights can be expressed as:

### v6

@@ Abstract
Large Language Models (LLMs) have fundamentally altered how we approach scaling in machine learning. However, these models pose substantial computational and memory challenges, primarily due to the reliance on matrix multiplication (MatMul) within their attention and feed-forward (FFN) layers. We demonstrate that MatMul operations can be eliminated from LLMs while maintaining strong performance, even at billion-parameter scales. Our MatMul-free models, tested on models up to 2.7B parameters, are comparable to state-of-the-art pre-trained Transformers, and the performance gap narrows as model size increases.
Our approach yields significant memory savings: a GPU-efficient implementation reduces memory consumption by up to 61% during training and over 10$\times$ during inference. When adapted for a multi-chip neuromorphic system, the model leverages asynchronous processing to achieve 4$\times$ higher throughput with 10$\times$ less energy than edge GPUs. These findings demonstrate a path toward dramatically simplified yet effective LLMs, advancing them toward brain-like efficiency and heralding a new generation of lightweight, high-performance language models. Our code implementation is available at https://github. com/ridgerchu/matmulfreellm.
1 Main
@@
Matrix Multiplication (MatMul) is the dominant operation in most neural networks, where dense layers involve vector-matrix multiplication (VMM), convolutions can be implemented as block-sparse VMMs with shared weights, and self-attention relies on matrix-matrix multiplication (MMM). The prevalence of MatMul is primarily due to Graphics Processing Units (GPUs) being optimized for MatMul operations. By leveraging Compute Unified Device Architecture (CUDA) and highly optimized linear algebra libraries such as cuBLAS, the MatMul operation can be efficiently parallelized and accelerated. This optimization was a key factor in the victory of AlexNet in the ILSVRC2012 competition and a historic marker for the rise of deep learning krizhevsky2012imagenet . AlexNet notably utilized GPUs to boost training speed beyond CPU capabilities, and as such, deep learning won the ‘hardware lottery’ hooker2021hardware . It also helped that both training and inference rely on MatMul.
@@
Despite its prevalence in deep learning, MatMul operations account for the dominant portion of computational expense, often consuming the majority of the execution time and memory access during both training and inference phases. Several works have replaced MatMul with simpler operations through two main strategies. The first strategy involves substituting MatMul with elementary operations, e.g., AdderNet replaces multiplication with signed addition in convolutional neural networks (CNNs) chen2020addernet . Given the focus on convolutions, AdderNet is intended for use in computer vision over language modeling. ShiftAddLLM enables re-parameterization of dense layers, though attention layers still rely on dynamic MatMul operations you2024shiftaddllm , where I/O limits will quickly overtake arithmetic savings as the sequence length grows.
@@
The second approach employs binary or ternary quantization, simplifying MatMul to operations where values are either flipped or zeroed out before accumulation. Quantization can be applied to either activations or weights: spiking neural networks (SNNs) use binarized activations maass1997networks ; eshraghian2023training ; zhu2023spikegpt , while binary and ternary neural networks (BNNs and TNNs) use quantized weights courbariaux2016binarized . Both methods can also be combined venkatesh2024squat ; eshraghian2022memristor .
@@
Recent advances in language modeling, like BitNet and Falcon-Edge wang2023bitnet ; ma20241bitllm ; tiionebitllms , demonstrate quantization’s scalability, replacing all dense layer weights with binary/ternary values to support up to 3B parameters. Despite replacing VMMs with accumulations in all dense layers, they retain the self-attention mechanism which relies on an expensive MMM. Dynamically computed matrices $Q$ (query) and $K$ (key) are multiplied to form the attention map. Since both $Q$ and $K$ matrices are dynamically computed from pre-activation values, achieving optimal hardware efficiency on GPUs requires custom optimizations, such as specialized kernels and advanced memory access patterns. Despite these efforts, such MatMul operations remain resource-intensive on GPUs, as they involve extensive data movement and synchronization which can significantly hinder computational throughput and efficiency dao2022flashattention . In our experiments, ternary quantization of the attention matrices in BitNet causes a significant drop in performance and failure to reach model convergence (see Fig. 4). This raises the question: is it possible to completely eliminate MatMul from LLMs?
@@
In this work, we develop the first scalable MatMul-free language model (Matmul-free LM) by using additive operations in dense layers and element-wise Hadamard products for self-attention-like functions. Specifically, ternary weights eliminate MatMul in dense layers, similar to BNNs. To remove MatMul from self-attention, we optimize the Gated Recurrent Unit (GRU) cho2014learning to rely solely on element-wise products and show that this model competes with state-of-the-art Transformers while eliminating all MatMul operations.
To fully exploit the efficiency potential of our architecture, we developed microcode-level optimization to process this model on a neuromorphic cluster using Intel’s Loihi 2 platform. The architecture of our MatMul-free LM naturally aligns with neuromorphic computing paradigms, allowing us to achieve remarkable efficiency gains. By mapping our model to Loihi 2’s mesh of asynchronous neurocores, we surpass human-readable throughput by approximately 8$\times$ at 4.2 W power consumption, representing a significant advancement over conventional hardware. For autoregressive generation, our model maintains consistent 59.4 tokens/second throughput at a highly efficient 70.8 mJ/token, far outperforming comparable Transformer models on embedded GPUs by at least 4$\times$ in throughput and 10$\times$ in energy efficiency, demonstrating how neuromorphic hardware can effectively harness the MatMul-free properties of our architecture. This implementation moves LLMs closer to brain-like efficiency and points towards a new generation of lightweight, high-performance language models.
2 Building the MatMul-free Language Model
Figure 1: Overview of the MatMul-free LM. Left: general architecture of proposed model. Middle-right: Algorithm mapping of a single block across neurocores on a single Loihi2 chip. Top-Right: Multi-chip Dataflow for autoregression and pre-fill. During autoregression, only one chip consumes non-negligible dynamic power dissipation due to the clock-free system. Bottom-right: The MatMul-free LM is deployed on the Hala Point system which consists of 1,152 Loihi 2 chips.
A closer examination of the Transformer architecture reveals two fundamental components that heavily rely on MatMul operations. The first is the dense layer, which not only transforms input hidden states to create Query, Key, and Value matrices for attention computation but also serves as the core structure in Feed-Forward Network (FFN) layers. Indeed, dense layers account for nearly all parameters in Transformer models. The second MatMul-dependent component is the attention mechanism itself, which performs MatMul operations directly on activations to compute attention scores.
The challenge, therefore, is to eliminate MatMul operations from both these structures while maintaining comparable performance. Following the Metaformer yu2022metaformer framework, we can decompose the Transformer architecture into two essential functions: token mixing and channel mixing. The token mixer operates on input sequences (implemented as attention in traditional Transformers), while the channel mixer processes information across embedding dimensions (implemented as FFN layers in Transformers).
Our solution addresses both components through a comprehensive architectural transformation. We replace the attention mechanism with an element-wise recurrent neural network (RNN) that provides similar token mixing capabilities but relies solely on element-wise operations. Simultaneously, we transform all dense layers to use ternary weights, effectively eliminating MatMul operations throughout the entire architecture. This approach maintains the core functional properties of Transformers while dramatically reducing computational complexity. This architectural transformation achieves a complete elimination of MatMul operations through carefully designed structural modifications, while preserving the essential mixing capabilities that make Transformers effective. Specific implementation details are provided in 6. The following sections detail how each component is optimized and demonstrate the effectiveness of this MatMul-free approach.
3 Scaling Analysis
Neural scaling laws posit that model error decreases as a power function of training set size and model size. Such projections become important as training becomes increasingly expensive with larger models. A widely adopted best practice in LLM training is to first test scalability with smaller models, where scaling laws begin to take effect scalinglawopenai ; chichinllascalinglaw ; yang2023baichuan . The GPT-4 technical report revealed that a prediction model just $1/10,000$ the size of the final model can still accurately forecast the full-sized model performance gpt4 .
4 Performance on Downstream Tasks
In line with benchmarking in BitNet, we evaluated the zero-shot performance of these models on a range of language tasks, including ARC-Easy arc , ARC-Challenge arc , Hellaswag hellaswag , Winogrande winoGrande , PIQA piqa , and OpenbookQA openbookqa . The results are shown in Tab. 1. All evaluations are performed using the LM evaluation harness eval-harness . The MatMul-free LM models achieve competitive performance compared to the Transformer++ baselines across all tasks, demonstrating its effectiveness in zero-shot learning despite the absence of MatMul operations, and the lower memory required from ternary weights. Notably, the 2.7B MatMul-free LM model outperforms its Transformer++ counterpart on ARC-Challenge and OpenbookQA, while maintaining comparable performance on the other tasks. As the model size increases, the performance gap between MatMul-free LM and Transformer++ narrows, which is consistent with the scaling law. These results highlight that MatMul-free architectures are capable of achieving strong zero-shot performance on a diverse set of language tasks, ranging from question answering and commonsense reasoning to physical understanding.
Models Size ARCe ARCc HS OQ PQ WGe Avg.
370M parameters with 15B training tokens, Layer=24, d=1024
Transformer++ 370M 45.0 24.0 34.3 29.2 64.0 49.9 41.1
MatMul-free RWKV-4 370M 44.7 22.8 31.6 27.8 63.0 50.3 40.0
Ours 370M 42.6 23.8 32.8 28.4 63.0 49.2 40.3
1.3B parameters with 100B training tokens, Layer=24, d=2048
Transformer++ 1.3B 54.1 27.1 49.3 32.4 70.3 54.9 48.0
MatMul-free RWKV-4 1.3B 52.4 25.6 45.1 31.0 68.2 50.5 45.5
Ours 1.3B 54.0 25.9 44.9 31.4 68.4 52.4 46.2
2.7B parameters with 100B training tokens, Layer=32, d=2560
Transformer++ 2.7B 59.7 27.4 54.2 34.4 72.5 56.2 50.7
Ours 2.7B 58.5 29.7 52.3 35.4 71.1 52.1 49.9
Figure 3: Performance comparison and analysis of different models and configurations. (a) and (b) show the training performance comparison between Vanilla BitLinear and Fused BitLinear in terms of time and memory consumption as a function of batch size. (c) compares the inference memory consumption and latency between MatMul-free LM and Transformer++ across various model sizes.
5 Deployment on Neuromorphic Hardware
5.1 Training Efficiency Optimization on GPU
Our experiments show that our fused operator benefits from larger batch sizes in terms of faster training speeds and reduced memory consumption. When the batch size is $2^{8}$, the training speed of the 1.3B parameter model improves from 1.52s to 1.21s per iteration, a 25.6% speedup over the vanilla implementation. Additionally, memory usage decreases from 82GB to 32GB, a 61.0% reduction. The performance of the fused implementation improves significantly with larger batch sizes, allowing more samples to be processed simultaneously and reducing the total number of iterations.
5.2 Inference Efficiency Optimization on GPU
Fig. 3(c) presents a comparison of GPU inference memory consumption and latency between the proposed MatMul-free LM and Transformer++ for various model sizes. In the MatMul-free LM, we employ BitBLAS bitblas for acceleration to further improve efficiency. The evaluation is conducted with a batch size of 1 and a sequence length of 2048. The MatMul-free LM consistently demonstrates lower memory usage and latency compared to Transformer++ across all model sizes. For a single layer, the MatMul-free LM requires only 0.12 GB of GPU memory and achieves a latency of 3.79 ms, while Transformer++ consumes 0.21 GB of memory and has a latency of 13.87 ms. As the model size increases, the memory and latency advantages of the MatMul-free LM become more pronounced. It is worth noting that for model sizes larger than 2.7B, the results are simulated using randomly initialized weights. For the largest model size of 13B parameters, the MatMul-free LM uses only 4.19 GB of GPU memory and has a latency of 695.48 ms, whereas Transformer++ requires 48.50 GB of memory and exhibits a latency of 3183.10 ms. These results highlight the efficiency gains achieved by the MatMul-free LM, making it a promising approach for large-scale language modeling tasks, particularly during inference.
5.3 Neuromorphic Computing with Intel Loihi 2
While GPUs excel at conventional deep learning workloads grounded in dense floating-point matrix multiplications, the MatMul-free LM we present is dominated by low-precision element-wise operations. The arithmetic intensity is so low that many CUDA cores remain idle during inference. Additionally, using ternary weights in matrix multiplications naturally induces unstructured sparsity of $\approx 35\%$ (for the 370M model) which cannot easily be accelerated on GPUs, resulting in redundant calculations and memory movement. Our MatMul-free LM maps naturally onto large-scale neuromorphic hardware that executes element-wise, recurrent, low-precision operations close to memory. Intel’s Loihi 2, for instance, offers microcode-programmable neuron dynamics and on-chip support for sparse low-precision arithmetic, making it an ideal substrate for the MatMul-free model davies_advancing_2021 . Processing is done locally within 120 fully asynchronous “neurocores” on each chip, with the option of connecting up to 1,152 chips together into a larger processing system, see Fig. 6 in Appendix C. Each neurocore stores weights and recurrent states in its local SRAM, which minimizes memory movement and consequently improves both energy efficiency and latency.
Table 2: Quantization results for the 370M model, using PyTorch on GPU. All weight matrices are ternary. W8: using 8-bit element-wise weights. A8/A16: using 8-bit or 16-bit activations.
ARCc ARCe HS OQ PQ WGe Avg.
Transformer++ 24.0% 45.0% 34.3% 29.2% 64.0% 49.9% 41.1%
Ours 23.8% 42.6% 32.8% 28.4% 63.0% 49.2% 40.3%
W8A8 28.3% 26.8% 26.1% 27.0% 52.7% 51.5% 35.4%
W8A16 23.0% 42.4% 32.4% 27.8% 63.0% 50.1% 39.8%
To accommodate Loihi 2’s low-precision fixed-point arithmetic, we studied the zero-shot accuracy of our 370M-parameter model under various quantization schemes (Table 2). Quantizing normalization parameters and activations to 8-bit weights and 16-bit activations (W8A16) preserves nearly the same performance as the original MatMul-free LM, whereas 8-bit activations (W8A8) reduce accuracy more noticeably, thus activations are kept in 16-bit precision. All quantized operators were implemented end-to-end in fixed-point arithmetic on Loihi 2, including a custom fixed-point RMSNorm layer (Appendix C.2).
Table 3: Throughput and energy efficiency for various transformer-based language models with at most 500M parameters running on the NVIDIA Jetson Orin Nano compared to our MatMul-free LM running on Intel’s Loihi 2. Bolded metrics are based on single-chip metrics and inter-chip communication results.
Throughput (tokens/sec) Efficiency (mJ/token)
Generate 500 1000 2000 4000 500 1000 2000 4000
Alireo-400M 14.3 14.9 15.0 14.7 723 719 751 853
Qwen2-500M 13.4 14.0 14.1 14.1 791 785 816 912
Ours (370M, 1-chip) 71.3 71.3 71.3 71.3 59 59 59 59
Ours (370M, system) 59.4 59.4 59.4 59.4 70.8 70.8 70.8 70.8
Prefill 500 1000 2000 4000 500 1000 2000 4000
Alireo-400M 849.4 1620 2858 3153 11.7 7.8 5.8 6.8
Qwen2-500M 627 909 1514 2639 17.9 13.9 9.5 6.7
Ours (370M, 1-chip) 13965 13965 13965 13965 2.8 2.8 2.8 2.8
Ours (370M, system) 11637 11637 11637 11637 3.4 3.4 3.4 3.4
The MatMul-free LM on Loihi 2 was characterized on an Oheo Gulch single-chip Loihi 2 system (N3C1 silicon) running NxKernel v0.2.0 and NxCore v2.5.8 (only accessible to Intel Neuromorphic Research Community members). The 1-chip case neglects inter-chip communication. Inter-chip communication causes a derived $\approx$20slowdown over the single chip case (Appendix C.4.1). Transformer LMs characterized on NVIDIA Jetson Orin Nano 8GB using MAXN power mode running Jetpack 6.2, TensorRT 10.3.0, CUDA 12.4. Energy values include CPU_GPU_CV, SOC, and IO components as reported by jtop 4.3.0. Performance results as of Jan 2025 and may not reflect all publicly available security updates. Results may vary.
Throughput is at least 4$\times$ higher for autoregressive generation and at least 3.6$\times$ higher for prefill. More details can be found in Appendix C.4.1. Table 3 contrasts the performance of the 370M MatMul-free model on Loihi 2 against Transformer baselines running on an NVIDIA Jetson Orin Nano, showing at least 1.7$\times$ less energy per token for prefill, and at least 10$\times$ less energy per token for auto-regressive generation. We compare against the 500M parameter Qwen2 model yang2024qwen2technicalreport , and also against the 400M parameter Alireo model alireo2024 which is representative of a small-scale Llama model llama3modelcard . We have included extended experimental results that compare against server-class GPUs (H100) in the Appendix to provide additional context (C.4), though omit them here due to the significant differences between memory and power draw.
6 Methods
This section first introduces MatMul-free BitLinear layers with ternary weights (${-1,0,+1}$), replacing multiplications with additions and negations to cut compute and memory while preserving expressiveness (Sec. 6.1). It then presents a hardware-efficient fused BitLinear implementation (Sec. 6.2) and assembles the full MatMul-free LM architecture (Sec. 6.3): a MatMul-free token mixer built on the Linear Gated Recurrent Unit (MLGRU) to capture sequence dependencies and a channel mixer that employs a BitLinear-based Gated Linear Unit, so the entire model relies solely on additions and element-wise products. Training details are summarized in Sec. 6.4.
@@ 6.1 MatMul-free Dense Layers with Ternary Weights
In a standard dense layer, the MatMul between the input $x\in\mathbb{R}^{d}$ and the weight matrix $W\in\mathbb{R}^{d\times m}$ can be expressed as:
@@
where $y\in\mathbb{R}^{m}$ is the output. To avoid using standard MatMul-based dense layers, we adopt BitNet to replace dense layers containing MatMuls with BitLinear modules, which use ternary weights to transform MatMul operations into pure addition operation with accumulation, i.e., ternary accumulation. When using ternary weights, the elements from the weight matrix $W$ are constrained to values from the set $\{-1,0,+1\}$. Let $\widetilde{\mathbf{W}}$ denote the ternary weight matrix. The MatMul with ternary weights can be expressed as: